OpenStudy (anonymous):

Determine whether Rolle's Theorem applies to the following function on the given interval. If so, find the point(s) that are guaranteed to exist by Rolle's Theorem. \[f(x)=\cos4x\] on\[[\frac{\pi}{ 8 },\frac{ 3\pi }{ 8}]\]

OpenStudy (anonymous):

check to see if \(\cos(\frac{4\pi}{8})=\cos(\frac{12\pi}{8})\)

OpenStudy (anonymous):

\[\cos(\frac{4\pi}{8})=\cos(\frac{\pi}{2})=0\] and \[\cos(\frac{12\pi}{8})=\cos(\frac{3\pi}{2})=0\] so you are good to go

OpenStudy (anonymous):

take the derivative, set it equal to zero, and solve for \(x\) to finish the question

OpenStudy (anonymous):

the derivative would be -4sin4x right?